# Quantum backflow in biased tight-binding systems

### v1

@@ Abstract
We study the phenomenon of quantum backflow in tight-binding systems with complex couplings, considering different boundary conditions and lattice sizes. Backflow is an intrinsically non-classical effect where the density flux associated with a particle described by the superposition of wave functions with, say, positive momentum, acquires negative values. We calculate the superposition of positive momentum states that gives rise to the strongest backflow in the systems. We also evaluate the bounds on the total amount of probability that flows in the opposite direction of the particle’s momentum.
@@ I Introduction
The phenomenon of backflow is a non-classical and non-linear effect that occurs when the probability current (or density flux) of a particle takes values opposite to its momentum. This phenomenon was first identified in 1969 by Allcock in the context of arrival times in quantum mechanics ALLCOCK1969253. Bracken and Melloy Bracken_1994 systematically formulated the problem in terms of the probability current of a normalized wave packet with positive momentum. The most significant outcome of their approach is that the total probability that flows in the direction opposite to the momentum is bounded by a universal constant $c_{\text{BM}}$. The exact value of the constant $c_{\text{BM}}$ is unknown, with the best numerical estimates obtained in Penz_2006; eveson, suggesting that in the infinite continuous case considered in those works, $c_{\text{BM}}\approx 0.0384517$; while, to the best of our knowledge, what occurs in discrete systems is not known. The phenomenon of backflow has been investigated in various extensions of the original setup Dissipative; Goussev1; Melloy1; Goussev2, including relativistic generalizations Relativistic, scattering conditions Scatering, and many-particle formulations many_particle. Other systems that exhibit backflow have also been identified, for example in Berry_2010; Dijk; Strange_2024. Continuing in this vein, it is natural to extend the study of this phenomenon to discrete versions of such systems, and explore how they differ from their continuous counterparts. In this work we focus on the nature of backflow in tight-binding chains with constant (complex) next nearest-neighbor couplings and onsite energies, which can be regarded as discrete analogues of a free-particle–type Hamiltonian  Timothy. The tight-binding model has traditionally been used to determine electronic and transport properties in solids with periodic potentials, such as insulators and semiconductors ashcroft2011solid. However, its versatility and simplicity have allowed it to be used to characterize a wide variety of analogous physical systems, ranging from waveguides waveguide1; waveguide2; waveguide3; waveguide4 to quantum random walks qrw; qrw1. We consider both periodic and infinite chains, which can be related to the general backflow setups described in ring1 and Bracken_1994. Interest in periodic systems arises from several characteristics, among them that the momentum eigenfunctions are normalizable and their spectrum is discrete, making the calculation of the flux bounds easier to a certain extent ring1; Strange_2012. Generalizations using periodic boundary conditions have been used to study backflow in a variety of situations dirac_1_2; DIBARI2023128831; TwoParticlesBarbier_2025; Ring2. Extensions to two dimensional systems Paccoia, where it turns out that the backflow is not bounded twodimensions, have also been considered.
In what follows we address two questions that naturally arise in regards to this phenomenon, the first of which, somewhat surprisingly, has been rarely explored (see, for example, Ring2). These are: How large can the backflowing probability density flux be? and second, along the lines of Bracken_1994: Is the total probability flowing backwards bounded? We consider tight-binding chains with complex couplings. Among other features ComplexTB, such models have attracted attention because, they break time-reversal symmetry while remaining Hermitian.
@@
The outline of the paper is as follows: In Section II, we introduce the systems we will consider and present the solutions of the corresponding Schrödinger equations. In Section III, we determine the momentum operator associated with the system, with the aim of identifying the conditions under which its eigenvalues are positive, thereby enabling the construction of superpositions of states with strictly positive momentum. Section IV is devoted to the continuity equations and to the physical role of the imaginary part of the coupling constant in these systems. In Section V, we calculate the corresponding fluxes and analyze the optimal bounds of the probability current, both for a simple superposition of two states with positive momentum, and then for general superpositions of such states. For completeness, in section VI, we undertake the “traditional” calculation of Bracken and Melloy Bracken_1994; ring1, which consists in the computation of the total probability flowing backwards at the origin during a time interval $T$, in order to estimate the constant $c_{\text{BM}}$. Finally, our conclusions are summarized in Section VII.
@@ II The models and their solutions
In this section we present the two configurations of the system under study and their corresponding solutions. This shall also help to fix the notation followed throughout the paper.
The tight-binding chain is described by the Hamiltonian (see figure fig.1):
@@
where $\tau$ is a coupling strength, which sets the energy scale of the problem, and $\epsilon$ is a dimensionless parameter whose physical meaning will be discussed in Section IV. We are considering constant onsite energies, so without loss of generality, we can set these to zero. The Hamiltonian is written in the site basis and $\mathbf{S},\mathbf{S}^{\dagger}$ are shift operators: if a state is written in the site representation $\psi_{k}(j)$, the actions of these operators are Fukui; kelley2001difference
@@
The eigenstates state of the time-independent Schrödinger equation
@@
where $\xi=\arctan(\epsilon)$ and $k$ represents the pseudo-momentum (c.f. Sec. III for the discussion of the physical momentum).
Figure 1: Segment of the chain, the lattice elements are separated by a distance $(\Delta x)$ and the nearest neighbor hopping is characterized by a complex number with imaginary part $\pm\epsilon$.
For the periodic chain with $N$ sites the time-independent Schrödinger equation, its solutions (up to a normalization constant), and the corresponding energies are essentially the same as for the infinite chain with $k(n)=\frac{2n\pi}{N}$, where $n$ is an integer so the eigenfunction $\psi_{k(n)}(j)$ satisfies the boundary condition $\psi_{k(n)}(0)=\psi_{k(n)}(N)$ .
III The momentum operator in tight-binding systems
In this section, we determine the momentum operator $\mathbf{p}$ of a tight-binding system in order to establish the values of the pseudo-momentum $k$ for which the eigenvalues of the momentum operator are positive. This insures that the states in which there is backflow are composed only of superpositions of eigenstates with positive momentum.
We can write the momentum operator using the Heisenberg equation JTaguena; Gu,
where $\mu$ is the mass of the particle and the position takes the values $(\Delta x)j$, with $j$ integer (see fig.1Si_2025; Timothy .
The time evolution of the position operator can be obtained using the Baker–Campbell–Hausdorff lemma sakurai2017modern:
Computing the commutators we obtain
and
$\mathbf{H},[\mathbf{H},\mathbf{r}]]=0.$ (9)
$\mathbf{r}(t)=\mathbf{r}+\frac{\mathbf{p}}{\mu}t.$ (10)
As expected, this is the position operator for rectilinear motion, since the Hamiltonian is a discretization of a one dimensional system where the particle moves freely. The momentum operator is
and its eigenvalue equation is
This agrees with the well-known relation between energy (c.f. eq. (5)) and the momentum eigenvalue $p$ for tight-binding systems JTaguena; Pedersen; Lee2018; Boykin
From eq. (12) it can be seen that for the infinite chain, eigenstates with positive momenta occur when the following condition
$k\in\left[-\xi,\;\pi-\xi\right]$ (14)
is satisfied.
For the periodic chain the eigenstates with positive momentum will be those where
$n\in\left\{\eta_{1},\eta_{1}+1,\dots,\eta_{2}-1,\eta_{2}\right\},$ (15)
where $\eta_{1}=\left\lceil-\dfrac{\xi}{2\pi}\,N\right\rceil$ and $\eta_{2}=\;\left\lfloor\dfrac{\pi-\xi}{2\pi}\,N\right\rfloor$. Here $\lfloor x\rfloor$ and $\lceil x\rceil$ denotes the usual floor and ceil functions.
IV Continuity equation
Given a wave function $\Psi(j,t)$, solution of the time dependent Schrödinger equation with Hamiltonian (1), the continuity equation for its probability density $|\Psi(j,t)|^{2}$ reads
@@ IV Continuity equation
where $J(j,t)$ is the flux from site j1j-1 to $j$, at time $t$, and is given by
$J(j,t)=\frac{\tau}{i\hbar}\left(\Psi^{*}(j-1,t)\Psi(j,t)-\Psi(j-1,t)\Psi^{*}(j,t)+i\epsilon(\Psi^{*}(j-1,t)\Psi(j,t)+\Psi(j-1,t)\Psi^{*}(j,t))\right).$ (17)
To address the physical meaning of the parameter $\epsilon$ we consider the continuous limit of eq. (16), which can be obtained by recalling that the position $x$ is measured in terms of the fundamental length $(\Delta x)$ through the relation $x=(\Delta x)j$ weiss1994aspects. Now, taking $(\Delta x)\rightarrow 0$, we can expand the wave function in Taylor series as
$\Psi((\Delta x)(j\pm 1),t)\sim\Psi(x,t)\pm(\Delta x)\frac{\partial}{\partial x}\Psi(x,t)+\frac{(\Delta x)^{2}}{2}\frac{\partial^{2}}{\partial x^{2}}\Psi(x,t)+\dots,$ (18)
Using eqs.(16) and (17), we obtain that the continuous expression of the flux is
@@
where the constants $\tau$ and $\epsilon$ are chosen as:
@@
consistently with the finite-difference approximation Timothy; Inui_2023.
The first terms on the RHS of eq.(19) are the "standard" probability flux of a quantum particle, while the last can be identified as a "drift" term, much like what happens in biased diffusion weiss1994aspects. Applying the same approximations to the Hamiltonian in eq.(1) yields the time dependent Schrödinger equation
@@
A physical realization of a quantum system with such a drift can be obtained, for example, by considering the Hamiltonian of a free particle confined to a ring in the $xy$ plane with radius $r_{0}$, centered at the origin, in the presence of a constant magnetic field of the form $\vec{B}=B_{0}\hat{z}$ GVugalter_2004. The Hamiltonian for a free particle in an external magnetic field reads
@@
where the operator $\mathbf{L}_{z}$ is defined as usual by $\mathbf{L}_{z}=-i\hbar\frac{\partial}{\partial\vartheta}.$ The term proportional to $B_{0}^{2}$, after expanding the square, produces an irrelevant constant energy offset which can be ignored, so the Hamiltonian eq. (25) reduces to
@@
This is the Hamiltonian of a rigid rotor in a magnetic field GVugalter_2004, which is analogous to the Hamiltonian in equation (21), and can be regarded as the continuous counterpart of the Hamiltonian in eq. (1) with periodic boundary conditions.
A different interpretation is reached recalling that appearance of the drift term in diffusive systems can be associated with a moving reference frame, in which case such term can be eliminated by means of a simple Galilean transformation. It is easy to see that with the transformation Rosen; Greenberger
@@
in the time dependent Schrodinger equation eq. (21), we get rid of the drift-like flux and recover the standard quantum flux. The drift is also affects in the energy and momentum, as shown in expressions (5) and (12); and has the effect of changing the range of pseudomomentum values $k$ that correspond to states with positive momentum, as shown in eq. (14).
@@ V Maximal backflow and bounds on probability density flow
In this section we construct superpositions of states with positive momenta to study the appearance of backflow and determine the cases that maximize the effect. First, we explore the phenomenon in a two state superposition, and then analyze the phenomenon in general superpositions involving all possible states with positive momentum.
@@ V.1 An illustrative example: The two-state backflow
We consider a wave function $\Psi(j,t)$ comprised by two eigenstates $\psi_{k(m_{1})}(j,t)$ and $\psi_{k(m_{2})}(j,t)$ with $m_{1},m_{2}$ belonging to the set of eq. (15) so $\Psi(j,t)$ is the simplest superposition of eigenstates that have positive momentum:
$\Psi(j,t)=\frac{\cos\left(\frac{\theta}{2}\right)e^{i\left(\frac{2\pi m_{1}j}{N}-\frac{E_{m_{1}}}{\hbar}t\right)}+\sin\left(\frac{\theta}{2}\right)e^{i\left(\frac{2\pi m_{2}j}{N}-\frac{E_{m_{2}}}{\hbar}t+\gamma\right)})}{\sqrt{N}},$ (28)
@@
Note that the term $\sin\left(\frac{(m_{1}+m_{2})\pi}{N}+\xi\right)$ always takes positive values for the current election of $m_{1}$ and $m_{2}$. This can be seen by assuming, without loss of generality, that $m_{1}<m_{2}$, then $m_{1}\leq\frac{m_{1}+m_{2}}{2}\leq m_{2}$, so $\frac{(m_{1}+m_{2})\pi}{N}$ corresponds to a point between the bounds of the set defined in eq. (15). Thus, the minimum of the flux for a given $j$ and $\gamma$ fixed, is reached at times $t$ such that the cosine factor in which $t$ appears is -1, that is
@@
Now we can adjust the parameter $\theta$ to obtain the minimum flux given the two states with positive momenta. First, we rewrite the last equation as
@@
Note that $a<\sqrt{b^{2}+c^{2}}$ as long as $m_{1}\neq m_{2}$, so the lowest flux attainable from the superposition of plane waves with positive momenta is, indeed, negative.

### v2

@@ Abstract
We study the phenomenon of quantum backflow in tight-binding systems with complex couplings, considering different boundary conditions and lattice sizes. Backflow is an intrinsically non-classical effect where the density flux associated with a particle described by the superposition of wave functions with, say, positive-momentum, acquires negative values. We calculate the superposition of positive-momentum states that gives rise to the strongest backflow in the system. We also evaluate the bounds on the total amount of probability that flows in the opposite direction of the particle’s momentum.
@@ I Introduction
The phenomenon of backflow is a non-classical and non-linear effect that occurs when the probability current (or density flux) of a particle takes values opposite to its momentum. This phenomenon was first identified in 1969 by Allcock in the context of arrival times in quantum mechanics Allcock (1969). Bracken and Melloy Bracken and Melloy (1994) systematically formulated the problem in terms of the probability current of a normalized wave packet with positive-momentum. The most significant outcome of their approach is that the total probability that flows in the direction opposite to the momentum is bounded by a universal constant $c_{\text{BM}}$. The exact value of the constant $c_{\text{BM}}$ is unknown, with the best numerical estimates obtained in Penz et al. (2005); Eveson et al. (2005), suggesting that in the infinite continuous case considered in those works, $c_{\text{BM}}\approx 0.0384517$.
@@
The phenomenon of backflow has been investigated in various extensions of the original setup Mousavi and Miret-Artés (2020); Goussev (2019); Melloy and Bracken (1998); Goussev (2020), including relativistic generalizations Ashfaque et al. (2019), scattering conditions Bostelmann et al. (2017), and many-particle formulations Barbier (2020). Backflow has also been reported in other systems, such as quasi-stable quantum systems van Dijk and Toyama (2019), Hermite wave packets Strange (2024), and in connection with retropropagation and negative kinetic energy Berry (2010). Continuing in this vein, it is natural to extend the study of this phenomenon to discrete versions of such systems, and to explore how they differ from their continuous counterparts. In this work we focus on the nature of backflow in tight-binding chains with constant (complex) next nearest-neighbor couplings and onsite energies, which can be regarded as discrete analogues of a free-particle–type Hamiltonian  Boykin and Klimeck (2004). The tight-binding model has traditionally been used to determine electronic and transport properties in solids with periodic potentials, such as insulators and semiconductors Ashcroft and Mermin (2011). However, its versatility and simplicity have allowed it to be used to characterize a wide variety of physical systems, ranging from waveguides Szameit and Nolte (2010); Longhi (2009); Chen et al. (2021); Lederer et al. (2008) to quantum random walks Bromberg et al. (2009); Peruzzo et al. (2010). We consider both periodic and infinite chains, which correspond to discrete versions of the general backflow setups of particles confined to a continuous infinite straight line Bracken and Melloy (1994) and ring respectively Goussev (2021); Goussev et al. (2024). Interest in periodic systems arises from several characteristics, among them that the momentum eigenfunctions are normalizable and their spectrum is discrete, making the calculation of the flux bounds easier to a certain extent Goussev (2021); Strange (2012). Other generalizations using periodic boundary conditions have been employed to study backflow in a variety of situations, such as relativistic spin-$1/2$ particles Su and Chen (2018), massless charged Dirac fermions Di Bari et al. (2023), and systems of two identical particles Barbier and Goussev (2025). Extensions to two-dimensional systems Paccoia et al. (2020), where it turns out that the backflow is not bounded Barbier et al. (2023), have also been considered.
Recently, complex couplings in one-dimensional tight-binding chains have attracted interest, among other features, for breaking time-reversal symmetry while the Hamiltonian remains Hermitian Rojas-Rojas et al. (2024). In the context of this work, from a transport perspective, their relevance lies in the fact that the imaginary part of the hopping parameter can be interpreted as a bias, analogous to that found in classical diffusive systems Zimborás et al. (2013); Lu et al. (2016); Liu and Zhou (2015). As will be explained in more detail in the paper, this results in an additional term in the expression for the probability density flux. Considering this, we address two questions that naturally arise regarding quantum backflow: How large can the backflowing probability density flux be? and second, along the lines of Bracken and Melloy (1994): Is the total probability flowing backwards bounded? Surprisingly, the first question has rarely been explored. Recently, Goussev Goussev et al. (2024) was among the first to delve into this subject, obtaining the optimal bounds of the probability density flux for a finite superposition of low-energy eigenstates of the Hamiltonian of a particle confined to a ring. In a similar vein, we construct wave packets composed exclusively of positive-momentum states that exhibit the most pronounced instantaneous negative flux for both infinite and periodic chains. In this respect, it is worth noting that tight-binding systems, unlike their continuous counterparts, exhibit a bounded energy spectrum as a consequence of lattice discretization, which implies that it is not necessary to restrict the number of admissible non-negative momentum eigenstates to avoid infinite energies Goussev et al. (2024). Moreover, the inclusion of a bias results in an enhancement of the flux amplitudes as its strength is increased, while simultaneously allowing controlled adjustments of the spectral window of suitable eigenstates. On the other hand, concerning the second question, we find that, regardless of the system topology, the total integrated backflow in the simple tight-binding systems we consider, exceeds that of their continuous counterparts in certain parameter regions. In this respect, both approaches highlight complementary but related aspects of quantum backflow: the first leads to a more pronounced but shorter-lived negative flux whereas the second yields an oscillatory flux characterized by negative-valued oscillations, albeit with a more moderate amplitude.
@@
The outline of the paper is as follows: In Section II, we introduce the systems we will consider and present the solution of the corresponding Schrödinger equation. Section III is devoted to the continuity equations and to the physical role of the imaginary part of the coupling constant in these systems. In Section IV, we determine the momentum operator associated with the system, with the aim of identifying the conditions under which its eigenvalues are positive, thereby enabling the construction of superpositions of states with strictly positive-momentum. In Section V, we calculate the corresponding fluxes and analyze the optimal bounds of the probability current, considering both simple superpositions of two positive-momentum states and general superpositions thereof. For comparison, in Section VI, we undertake the “traditional” calculation of Bracken and Melloy Bracken and Melloy (1994); Goussev (2021), which consists in the computation of the total probability flowing backwards at the origin during a time interval $T$, in order to estimate the constant $c_{\text{BM}}$. Finally, our conclusions are summarized in Section VII.
@@ II The models and their solutions
In this section we describe the two configurations of the tight-binding model under study, and their corresponding solutions. This shall also help to fix the notation followed throughout the paper.
The tight-binding chain is described by the Hamiltonian (see fig.1):
@@
where $\tau$ is a coupling strength, which sets the energy scale of the problem, and $\epsilon$ is a dimensionless parameter whose physical meaning will be discussed in Section III. We are considering constant onsite energies, so without loss of generality, we can set these to zero. The Hamiltonian in discrete configuration space is expressed using the shift operators $\mathbf{S}$ and $\mathbf{S}^{\dagger}$. For a state in the site representation, $\psi_{k}(j)$, the actions of these operators are Fukui (2020); Kelley and Peterson (2001)
@@
Here we observe that $\mathbf{H}\neq\mathbf{H}^{*}$, indicating the absence of time-reversal symmetry, however, as $\mathbf{H}=\mathbf{H}^{\dagger}$, the Hamiltonian remains Hermitian. The eigenstates of the time-independent Schrödinger equation
@@
where $\xi=\arctan(\epsilon)$ and $k$ represents the pseudomomentum. For the purposes of this work, the pseudomomentum serves as a quantum number Streib (2021) taking values in the interval $\pi,\pi]$ for the infinite chain. This quantity is not to be confused with the physical momentum operator $\mathbf{p}$ nor its eigenvalues, discussed in Sec. IV.
Figure 1: Segment of the chain, the lattice elements are separated by a distance $(\Delta x)$ and the nearest-neighbor hopping is characterized by a complex number with imaginary part $\pm\epsilon$.
When we impose periodic boundary conditions on the chain of $N$ sites ($\psi_{k}(0)=\psi_{k}(N)$), the pseudomomentum takes discrete values
$k=\frac{2n\pi}{N},$
where $n$ is an integer ranging from 0 to $N$.
III Continuity equation
As mentioned in the previous section, we now provide a detailed discussion of the physical significance and implications of the "bias" parameter $\epsilon$. Let us consider a wave function $\Psi(j,t)$, solution to the time-dependent Schrödinger equation with Hamiltonian (1). The continuity equation for its probability density $|\Psi(j,t)|^{2}$ reads
@@ III Continuity equation
where $J(j,t)$ is the flux from the site j1j-1 to $j$, at time $t$, and is given by
$J(j,t)=\frac{\tau}{i\hbar}\Big(\Psi^{*}(j-1,t)\Psi(j,t)-\Psi(j-1,t)\Psi^{*}(j,t)+i\epsilon\big(\Psi^{*}(j-1,t)\Psi(j,t)+\Psi(j-1,t)\Psi^{*}(j,t)\big)\Big).$ (7)
Now, we consider the continuous limit of eq.(6), which can be obtained by recalling that the position $x$ is measured in terms of the fundamental length $(\Delta x)$ through the relation $x=(\Delta x)j$ Weiss (1994). Then, taking $(\Delta x)\rightarrow 0$, we can expand the wave function in a Taylor series as
$\Psi((\Delta x)(j\pm 1),t)\simeq\Psi(x,t)\pm(\Delta x)\frac{\partial}{\partial x}\Psi(x,t)+\frac{(\Delta x)^{2}}{2}\frac{\partial^{2}}{\partial x^{2}}\Psi(x,t)+\dots,$ (8)
Using eqs.(6) and (7), we obtain that the continuous expression for the flux is
@@
where the constants $\tau$ and $\epsilon$ are given by:
@@
consistently with the finite-difference approximation Boykin and Klimeck (2004); Inui (2023).
The first terms on the RHS of eq.(9) are the "standard" probability flux of a quantum particle, while the last can be identified as a "drift" term, much like what happens in biased diffusion Weiss (1994). Applying the same approximations to the Hamiltonian in eq.(1) yields the time-dependent Schrödinger equation
@@
A physical realization of a quantum system with such a drift can be obtained, for example, by considering the Hamiltonian of a free particle confined to a ring in the $xy$ plane with radius $r_{0}$, centered at the origin, in the presence of a constant magnetic field of the form $\vec{B}=B_{0}\hat{z}$ Vugalter et al. (2003). The Hamiltonian for a free particle in an external magnetic field reads
@@
where the operator $\mathbf{L}_{z}$ is defined as usual by $\mathbf{L}_{z}=-i\hbar\frac{\partial}{\partial\vartheta}.$
The term proportional to $B_{0}^{2}$, after expanding the square, produces an irrelevant constant energy offset which can be ignored, so the Hamiltonian eq.(15) reduces to
@@
This is the Hamiltonian of a rigid rotor in a magnetic field Vugalter et al. (2003), which is analogous to the Hamiltonian in equation (11), and can be regarded as the continuous counterpart of the Hamiltonian in eq. (1) with periodic boundary conditions.
A different interpretation is reached recalling that the appearance of the drift term in diffusive systems is associated with a moving reference frame, in which case such term can be eliminated by means of a simple Galilean transformation. It is easy to see that with the transformation Rosen (1972); Greenberger (2001)
@@
in the time-dependent Schrödinger equation eq.(11), we get rid of the drift-like flux and recover the standard quantum flux.
The drift also has the effect of changing the range of pseudomomentum values $k$ that correspond to states with positive-momentum, as we show in the next section.
IV The momentum operator in tight-binding systems
Before starting our analysis of quantum backflow in this system, it is necessary to determine the momentum operator $\mathbf{p}$ of a biased tight-binding chain in order to establish the values of the pseudomomentum $k$ for which its eigenvalues are non-negative.
We can write the momentum operator using the Heisenberg equation Cruz et al. (1999); Gu et al. (2013),
where $\mu$ is the mass of the particle and the position takes the values $(\Delta x)j$, with $j$ integer (see fig.1Si and Zhang (2025); Boykin and Klimeck (2004) .
The time evolution of the position operator can be obtained using the Baker–Campbell–Hausdorff lemma Sakurai and Napolitano (2017):
Computing the commutators yields
and
$\mathbf{H},[\mathbf{H},\mathbf{r}]]=0.$ (21)
$\mathbf{r}(t)=\mathbf{r}+\frac{\mathbf{p}}{\mu}t.$ (22)
As expected, this is the position operator for rectilinear motion, since the Hamiltonian is a discretization of a one-dimensional system where the particle moves freely. The momentum operator is
and its eigenvalue equation is
This agrees with the well-known relation between energy (c.f. eq. (5)) and the momentum eigenvalue for tight-binding systems Cruz et al. (1999); Pedersen et al. (2001); Lee et al. (2018); Boykin and Vogl (2001)
From eq. (24), it can be seen that for the infinite chain, eigenstates with positive-momenta occur when the following condition
$k\in\left[-\xi,\;\pi-\xi\right]$ (26)
is satisfied.
For the periodic chain; the eigenstates with positive-momentum will be those for which
$n\in\left\{\eta_{1},\eta_{1}+1,\dots,\eta_{2}-1,\eta_{2}\right\},$ (27)
where $\eta_{1}=\left\lceil-\dfrac{\xi}{2\pi}\,N\right\rceil$ and $\eta_{2}=\;\left\lfloor\dfrac{\pi-\xi}{2\pi}\,N\right\rfloor$. Here, $\lfloor x\rfloor$ and $\lceil x\rceil$ denote the usual floor and ceil functions.
@@ V Maximal backflow and bounds on probability density flow
We start our analysis of quantum backflow by obtaining the states that maximize the amplitude of the negative flux. First, we illustrate the key idea using a wave function composed of two eigenstates, and then extend the approach to wave packets formed from all positive-momentum eigenstates in both the periodic and infinite cases.
@@ V.1 An illustrative example: The two-state backflow
We consider a wave function $\Psi(j,t)$ comprised of two eigenstates $\psi_{k(m_{1})}(j,t)$ and $\psi_{k(m_{2})}(j,t)$ with $m_{1},m_{2}$ belonging to the set of eq. (27) so $\Psi(j,t)$ is the simplest superposition of eigenstates that have positive-momentum:
$\Psi(j,t)=\frac{\cos\left(\frac{\theta}{2}\right)e^{i\left(\frac{2\pi m_{1}j}{N}-\frac{E_{m_{1}}}{\hbar}t\right)}+\sin\left(\frac{\theta}{2}\right)e^{i\left(\frac{2\pi m_{2}j}{N}-\frac{E_{m_{2}}}{\hbar}t+\gamma\right)}}{\sqrt{N}},$ (28)
@@
Note that the term $\sin\left(\frac{(m_{1}+m_{2})\pi}{N}+\xi\right)$ always takes positive values for the current choice of $m_{1}$ and $m_{2}$. This can be seen by assuming, without loss of generality, that $m_{1}<m_{2}$, then $m_{1}\leq\frac{m_{1}+m_{2}}{2}\leq m_{2}$, so $\frac{(m_{1}+m_{2})\pi}{N}$ corresponds to a point between the bounds of the set defined in eq. (27). Thus, the minimum of the flux for a given $j$ and $\gamma$ fixed, is reached at times $t$ such that the cosine factor in which $t$ appears is -1, that is
@@
Now we can adjust the parameter $\theta$ to obtain the minimum flux given the two states with positive-momenta. First, we rewrite the last equation as
@@
Note that $a<\sqrt{b^{2}+c^{2}}$ as long as $m_{1}\neq m_{2}$, so the lowest flux attainable from the superposition of plane waves with positive-momenta is, indeed, negative.